Solve by factoring: (x^2/4) – 3 – 5x = – 3
Answer by user_dude2008(1825)

(x^2/4) – 3 – 5x = – 3


x^2 – 12 – 20x = – 12


x^2 – 20x = 0

x(x – 20) = 0

Ans: x = 0 or x = 20